OpenStudy (anonymous):

need help with a calculus problem :/ Compute f '(a) algebraically for the given value of a: f(x) = x3 + 2x; a = 9

OpenStudy (anonymous):

Do you mean using the limit definition?

OpenStudy (anonymous):

its x^3

OpenStudy (anonymous):

f'(x) = 3x^2 + 2. Plug in 9. 3(9)^2 + 2 = 245

OpenStudy (anonymous):

f'(x) = 3x^2 + 2 f'(9) = 3(9)^2 + 2 f'(9) = 245

OpenStudy (anonymous):

thats what i got thanks guys